OpenStudy (anonymous):

how do a verify the expression sec^2 = 2tanx + 2tan^3x

OpenStudy (anonymous):

\[\sec^2x=2tanx+2\tan^2x\] definition, Sec^2x=Tan^2x+1 \[Tan^2x+1=2tanx+2\tan^3x\] Do it.

OpenStudy (raden):

proving or solving that equation ?

OpenStudy (anonymous):

Do you have to use one side independently?

OpenStudy (anonymous):

I'll assume no, OK?

OpenStudy (anonymous):

\[Tan^2x+1=2Tanx+2Tan^3x\]

OpenStudy (anonymous):

this is it, sorry again.

OpenStudy (anonymous):

\[Tan^2x+1=2Tanx(Tan^2+1)\] 2tanx has to be equal one. this is not an identity, it's an equation to solve

OpenStudy (anonymous):

\[x= Tan ^{-1}(1/2)\]

OpenStudy (anonymous):

Did I go too fast?

OpenStudy (anonymous):

Can someone help me with my literature question? http://openstudy.com/study#/updates/529a99bfe4b0a454724fff80

OpenStudy (anonymous):

\[Sec^2x=Tan^2x+1~~~~~~~~~~So,\]\[Tan^2x+1=2Tanx+2Tan^3x\]\[Tan^2x+1=2Tanx~(1+Tan^2x)~~~~~~can cel~~out~~(Tan^2x+1)\]\[2Tanx=1~~~~~~->~~~~~~~~~~~~Tan ^{-1}(1/2)\]

OpenStudy (anonymous):

that's it basically. Again, this is NOT an IDENTITY!

OpenStudy (raden):

the degrees of LHS is 2 while RHS is 3, obviously this cant be an identity :)

OpenStudy (anonymous):

This is a rate of change problem so I'm not sure if I have to prove it or solve it. The problem says to verify that the trigonometric expression for the rate of change is also 2 tanx + 2 tan^3x. the most confusing part about this problem is that I can't tell whether it's an identity or not.

OpenStudy (anonymous):

Right, it's not an identity, did you look at my work? I showed that it's not an identity. You can also check by plugging in values for x.

OpenStudy (anonymous):

Look, let's take 90. x=90 and Tan=Sin/Cos we can plug this into \[Tan^2x+1=2Tanx+2Tan^3x\] this is undefined b/c Co90=0 Do you understand this?